OpenStudy (anonymous):

lim n tends to infinity sum from k=0 to n (nCk)/ (n^k)(k+3) where nCk is n!/(k!(n-k)!) How do you express this sum as an integration?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[\frac{nCk}{ n^k (k+3)}\]?

OpenStudy (slaaibak):

Anyway you can give this more visually? I'm not 100% how the question looks like

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[\frac{nCk}{n^k}(k+3)\]

OpenStudy (anonymous):

k+3 is in the denominator

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

easier way is start with the expansion (1+x)^n

OpenStudy (anonymous):

(1+x)^n = nC0 + nC1x + nC2 x^2 + .... nCk x^k + ... nCn x^n

OpenStudy (slaaibak):

\[\lim_{n \rightarrow \infty} \sum_{k=0}^{n} nCr (n^k)/(k+3) \] ?

OpenStudy (anonymous):

(1+n)^x = xC0 + xC1 n + xC2 n^2 + .... +xCk n^k + .... x Cx n^x

OpenStudy (anonymous):

n^k and k+3 are in the denominator numerator is only nCk

OpenStudy (anonymous):

well I know what the solution to the integral is, do you want to know that? maybe it would help you in deciding what the integral should look like. I'm not sure myself how to write the integral yet

OpenStudy (anonymous):

no i want what the integral is the answer is not really important...